if a boys age and his fathers age amount together to 24 years. fourth part of the product of their ages exceeds the boys age by 9 years.find how old they are.
Answer by mananth(16949)

boy's age =x
father's age = y
x+y=24
(xy)/4 -x=9
multiply equation by 4
xy-4x=36
x(y-4)=36
x=36/(y-4)
Now x+y = 24
substitute x in the equation
36/(y-4) +y = 24
multiply equation by (y-4)
36+y(y-4)=24(y-4)
36+y^2-4y=24y-96
y^2-4y-24y+36+96=0
y^2-28y+132=0
y^2-22y-6y+132=0
y(y-22)-6(y-22)=0
(y-22)(y-6)=0
y=22 OR 6
Father's age = 22
Son's age = 2 (because x+y=24)